Skysong Inc. issues $2,084,300 of 10% bonds due in 13 years with interest payable at year-end. The current market rate of interest for bonds of similar risk is 11%. Click here to view factor tables What amount will Skysong receive when it issues the bonds? (Round factor values to 5 decimal places, e.g. 1.25124 and final answers to 0 decimal places, e.g. 458,581.)
Answer #1

Formula : Value of bond = Coupon * PVIFA(market rate , n) + Par value * PVIF(market rate, n)

Value of bond = (2084300*10%) * PVIFA(11% , 13) + 2084300 * PVIF(11%, 13)

= 208430 * 6.750 + 2084300 * 0.258 = $1944651.90

Skysong Inc. will receive $1944651.90 on issue of the bonds.